OpenStudy (anonymous):

Subtracting Mixed Number 2 4/7- 1 3/5

OpenStudy (anonymous):

first convert these mixed numbers into improper fractions then do subtraction by LCM method

OpenStudy (karatechopper):

2 4/7- 1 3/5 Improper fraction 18/7 - 8/5 Find common denominator..35 18/7=?/35 =90/35 -8/5=?/35 -56/35 90/35-56/35 Simplify 34/35